OpenStudy (nuccioreggie):

Jason has two bags with 6 tiles each. The tiles in each bag are shown below: Make 6 squares. The squares are numbered sequentially from 1 to 6. Without looking, Jason draws a tile from the first bag and then a tile from the second bag. What is the probability of Jason drawing an even tile from the first bag and an even tile from the second bag? 6 over 12 9 over 12 6 over 36 9 over 36

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

I got A

OpenStudy (michele_laino):

the probability to draw an event tile from a single bag, is: p= favorable outcomes/ possible outcomes= 3/6

OpenStudy (michele_laino):

so, the requested probability, is the product of the two probabilities, to draw an even tile from each bag: p=p_1*p_2= (3/6) * (3/6)=...

OpenStudy (nuccioreggie):

6/12

OpenStudy (michele_laino):

are you sure? they are independent events

OpenStudy (nuccioreggie):

What do u think the answer choice is?

OpenStudy (michele_laino):

hint: